OpenStudy (anonymous):

Light with a wavelength of 5.0 · 10-7 m strikes a surface that requires 2.0 ev to eject an electron. Calculate the energy, in electron volts, of one incident photon

OpenStudy (sidsiddhartha):

energy=hc/lamda=(6.625*10^-34*3.8*10^8)/(5*10^-7)=3.97*10^-19j=2.484 ev